A professor is interested in the average length of books in her library. She has divided her books into a few different categories: 235 books on mathematics, 290 books on sports, and 166 books on interior design. Rather than examining all the books, she plans to use a stratified sample of 50 books. How many of the sports books should she choose? Enter a whole number.
Answer #1

Total number of books = 235 + 290 + 166 = 691

Proportion of sample to the population = 50/691

The proportion shall be maintained in all strata of the sample. So, the number of samples to be taken from sports books

= Number of sports books x 50/691

= 290 x 50/691

= 21